Find the time required for an investment of 1 dollars to grow to 6700 dollars at an interest rate of 7.5 percent per year, compounded quarterly.
Answer #1

We use the formula:
A=P(1+r/400)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

6700=1(1+7.5/400)^4n

6700=(1.01875)^4n

Taking log on both sides;

log 6700=4n*log (1.01875)

Hence n=1/4[log 6700/log (1.01875)]

which is equal to

=118.56 years(Approx).